Symmetry and numerical analysis of the flow of a thin non-Newtonian fluid on an inclined plane

Serge N. Neossi N. University of the Witwatersrand, South Africa

This talk considers the evolution equation, its symmetries , reductions to Ordinary Differential Equations and the numerical study of the development of the flow of a thin non-Newtonian liquid film on an inclined plane and investigates the effect of gravity, surface tension and Marangoni on the film height. Examples of non-Newtonian fluids are polymer solutions, soap, blood and paint.